Day shift 24 Nov. 1999

Nickolas Solomey (solomey@hep.uchicago.edu)
Wed, 24 Nov 1999 15:29:15 -0600

Shift Summary for owl shift, 24/Nov/99

On Shift: Nick and Roland

Hours of beam delivered at all intensities: 2.7
Hours of beam delivered at nominal intensity: 2.5
Hours of E799 data (special runs not included): 2.5
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: Ice ball in accelerator, warming up, plus
some LCW problems, all accelerator related problems.

========================================================================
Runs ended this shift: 15040, 15041, 15042

Tapes brought to FCC this shift: UPx178-180 (for x=A,B,C) and UPS060,
and UPX012.
========================================================================
Accesses during shift: Leo and Jason to change vv' timing, but in the
end they did nothing, then the beam vanished
and there were a lot plus a search and secure
party after the interlocks were dropped, see
description below.
========================================================================
Overview of shift: small amount of data taking the rest just
problem solving.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15040 07:00 Change of shift, good data taking in progress.
322 07:28 Acceleartor missed a spill.
362 Accelerator missed a spill.
08:45 Onsplit requested UPH032 mounted on ktev4.
396 LCW problems beam down, no estimate.
401 Run ended.
09:10 Updated DC t0 constants using file dct015040.
15041 Run started all was fine so it was ended to
change tapes.
15042 09:45 Run started with new tapes.
09:50-10:05 Jason and Leo made access to change vv' timing
cables, but they did not change anything
because it was changed so much they did not
recognize where it was.
20 10:15 Beam returns.
36 10:37 MCR calls; they will reduce intensity while
tuning for about 1/2 hour.
Beam went away, C1 cryo problems inhibiting
TeV ramp.
11:15 MCR called, they are making a MI access until
13:30.
65 Stopped run, to start cosmic run.
15043 11:20 Started cosmic run.
11:30 Valeri, Julie and Jason go on access.
12:00 Julie, Jason and Valeri come out of access.
12:02 MCR called, there is an ice ball in G2 they
have to warm it up to remove it and then
they will be ready to run the Tevatron
again in 4-8 hours, an estimate.
13:45 Rick Kessler called, he noticed were were
running cosmics and wants to work on
writing a KAMI file, which after hearing
the estimate of the time we will be down
he decided this was a good time to upgrade
the software.
14:10 Jason and Julie returned for another access, they
noticed that our interlocks were dropped. Called
MCR to inform and ask them if they know when.
The key tree was locked so no one else had
tried to go in since the previous access.
14:20 MCR came over, looked at the interlock electronics
rack and said it was our main door that dropped
the interlock.
15:05 Julie and Jason kicked out, Search and Secure
performed.
15:25 Nick and Bob T. made access to inspect RF cable
path to see if it is possible to adjust timing
from upstairs.
15:30 Shift ends, hope the next shifts and holiday
gets lots of beam.

=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=
Dr. Nickolas Solomey, Research Scientist, Enrico Fermi Institute
The University of Chicago, 5640 S. Ellis Ave, Chicago, Ill. 60637 USA
Phone: (773) 702-3278, fax: x-6317, http://hep.uchicago.edu/~solomey